# Revisiting the Berkeley Admissions data: Statistical Tests for Causal Hypotheses

### v1

@@ D.1 Sharpness of IV inequalities
For $M\in\mathbb{M}_{\text{IV}}$, the response-function parameterization yields a counterfactually equivalent SCM (Forré and Mooij, 2024, Section 8.2, 8.4) $\tilde{M}=(V,\tilde{W},\tilde{\mathcal{X}},\tilde{f},\tilde{P})$, where $V=\left\{Z,X,Y\right\},\tilde{W}=\left\{R,U_{Z}\right\},\tilde{\mathcal{X}}=\mathcal{X}_{V}\times\mathcal{X}_{\tilde{W}},\tilde{f}={\left({\tilde{f}_{Z},\tilde{f}_{X},\tilde{f}_{Y}}\right)}$ where we define $\mathcal{X}_{R},\tilde{f},\tilde{P}$ through the function $\Phi:\mathcal{X}_{W}\mapsto\mathcal{X}_{\tilde{W}}$ where
@@
$\mathbb{E}=\left\{\delta_{x,y|0}+\delta_{x^{\prime},y^{\prime}|1}:x,x^{\prime}\in{\cal X};y,y^{\prime}\in{\cal Y}:x\neq x^{\prime}\right\}\cup\left\{\delta_{x,y|0}+\delta_{x,y|1}:x\in{\cal X},y\in{\cal Y}\right\}$
@@
For a point to be an extreme point of $\mathcal{K}_{\text{IV}}$, it has to satisfy the above $26n+$ 2 constraints and additionally, at least $n_{X}n_{Y}n_{Z}=4n$ of them must be active and linearly independent. The normalization constraints have to be satisfied and therefore, active, which leaves us with enumerating all feasible points for which at least $24n-$ 2 of the $6n$ inequalities are active and linearly independent.
@@
We first show that all points in $\mathbb{E}$ are extreme points of $\mathcal{K}_{\text{IV}}$. Choose $x,x^{\prime}\in{\cal X}$ with $x\neq x^{\prime}$ and choose $y,y^{\prime}\in{\cal Y}$. It can be verified that $\delta_{x,y|0}+\delta_{x^{\prime},y^{\prime}|1}$ satisfies the IV inequalities with $2$ out of the $2n$ IV inequalities being active. Further, $24n-$ 2 non-negativity constraints are active. All active constraints are linearly independent. Choose $x\in{\cal X},y\in{\cal Y}$. It can be verified that $\delta_{x,y|0}+\delta_{x,y|1}$ satisfies the IV inequalities with $2$ out of the $2n$ IV inequalities being active. Further, $24n-$ 2 non-negativity constraints are active. All active constraints are linearly independent.
@@
Finally, we check whether $\mathbb{E}$ exhausts all extreme points. Pick a feasible point $b\in\mathbb{R}^{n_{X}n_{Y}n_{Z}}$. We will refer to the indices $i$ with $b_{i}\neq 0$ ($b_{i}=0$) as the “non-zero (zero) entries of $b$”, and to the indices $j$ in an active IV inequality constraint $a^{\top}b=1$ with $a_{j}\neq 0$ as the “active entries of the IV inequality”, and say that two such entries overlap if $i=j$. We also call the set of entries of $b$ corresponding to $K(X,Y|Z=z)$ a “stratum”. We proceed case by case.
1. More than two IV inequalities are active in $b$. This yields a contradiction with the normalization constraints.
@@
2. Exactly two IV inequalities are active in $b$. Then $b$ must contain at least $44n-$ 4 zero entries, of which at most two can overlap with the active IV inequality entries. a) If there are no overlaps, then $b$ must be 00 in all other entries, but then the active constraints would not be linearly independent, since the sum of the active normalization constraints equals equals the sum of active IV constraints and the active nonnegativity constraints. b) If at least one of these overlaps, then $b_{j}=1$ with $j$ the other active entry in the active IV inequality with the overlap. This implies $12n-$ 1 zeroes in the stratum of $j$, so in total we have identified $2n$ zeroes in $b$. We need $42n-$ 4 more active constraints. Consider the other IV inequality. Since $b_{j}=1$, the active entry corresponding to that stratum must be a zero entry of $b$. Hence, $b_{k}=1$ for $k$ the other active entry in this IV inequality. Thus this gives us the extreme points already identified.
3. Exactly one IV inequality is active in $b$. Then we need to pick $34n-$ 3 zero entries corresponding to active nonnegativity constraints. a) Suppose none of these entries overlaps. Then we have $24n-$ 2 locations for these zeros. We can pick one nonzero entry that’s not an active IV inequality entry. The normalization constraint implies $b_{j}=1$ for one of the two active IV inequality entries, and hence $b_{k}=0$ for the other one. Contradiction. b) Suppose exactly one of these entries overlaps. We need $22n-$ 2 more active constraints. Only two possible non-zero entries for $b$ remain. In order for the active constraints to span the entire vector space, at least one of these should be a zero entry. This can be checked by observing that if the two possible non-zero entries of $b$ are positive, then the $4n$ active constraints cannot express the vector that is $1$ in all entries except the two positive entries where it is 00. Therefore, if at least one of the two possible non-zero entries of $b$ is zero, then $b_{j}=1$ for the other entry, which activates another IV inequality. Contradiction. c) If there are two overlapping zeroes the IV inequality couldn’t be active.
@@
4. Zero IV inequalities are active in $b$. Then $b$ must contain at least $24n-$ 2 zero entries. Because of the normalization constraints, we need one non-zero entry in both strata, and it must actually be a $1$. This will activate at least one IV inequality. Contradiction.
@@ D.3 Equivalence of Statistical Tests
Note that ${\cal P}_{\text{IV}+}=\left\{P(Z):\forall z,P(Z=z)>0\right\}\otimes\mathcal{K}_{\text{IV}}$ since assuming positivity, (3) is identical to (4). Further, ${\cal P}_{\text{cf-notion}}=\left\{P_{M}(Z):M\in H^{0}_{\text{cf-notion}+}\right\}\otimes{\cal K}_{\text{cf-notion}}$ since assuming positivity, $\mathbb{M}_{\text{cf}+}=\mathbb{M}_{\text{cf}}$ and $P_{M}{\left({X,Y\mid\text{do}(Z)}\right)}=P_{M}{\left({X,Y\mid Z}\right)}$ for $M\in\mathbb{M}_{\text{cf}+}$. Since the first factors are identical, Theorem 18 follows from Lemma 26. ∎
@@ E.1.1 Counterfactual Fairness and Demographic Parity
We restate the counterfactual notion of fairness from Kusner et al. (2017) for the Berkeley example below.
Definition 27 (Counterfactual Fairness (Kusner et al., 2017)).
@@
The counterfactual fairness notion of Kusner et al. (2017) implies demographic parity for the Berkeley example without allowing for confounding.
@@
Note that demographic parity falls prey to Simpson’s paradox in the Berkeley example. The above result shows that a valid test for demographic parity is a valid test for Kusner et al. (2017)’s counterfactual fairness notion for the assumed model class, $\mathbb{M}_{\text{no-cf}}$.
@@ E.1.2 Path-dependent Counterfactual Fairness
We next show that testing the path-dependent counterfactual fairness notion given in the appendix of Kusner et al. (2017) coincides with a conditional independence test $A\perp\!\!\!\perp S\mid D$.
Definition 29 (Path-dependent Counterfactual Fairness (Kusner et al., 2017)).
@@ E.2 With Confounding
In this section, we compare the statistical tests that result from the NDE that the notions of Nabi and Shpitser (2018) and Chiappa (2019) are based on, and Kusner et al. (2017)’s counterfactual fairness and path-dependent counterfactual fairness notions, when confounding is allowed.
@@ E.2.1 NDE
With confounding between the mediator and the outcome, Kaufman et al. (2005) obtain bounds on the NDE for the all-binary variable case by the linear programming approach of Balke and Pearl (1997). The resulting bounds are implied by the IV inequalities but not equivalent to them, which gives us a strictly weaker test than the IV inequalities. For completeness, we present the bounds below. The lower and upper bounds for $NDE(A;0\rightarrow 1)$ are
@@
E.2.2 Counterfactual Fairness (Kusner et al., 2017)
@@
E.2.3 Path-dependent Counterfactual Fairness (Kusner et al., 2017)
@@ Appendix F Additional Results for Bayesian Testing Procedure: Cum-laude Dataset and Prior Sensitivity
We consider the dataset from Bol (2023) that contains data from $5239$ PhD students in the Netherlands studying at a large Dutch university from 2011-2021. Bol (2023) observed a bias in the percentage of ‘cum-laude’ distinctions awarded to male PhD students ($6.57\%$) versus female PhD students ($3.68\%$).
@@
As reported by Bol (2023), unlike the Berkeley dataset, the bias among female and male cum-laude award rates does not vanish when conditioned on department. Therefore, with the assumption of no confounding between the academic field choice and the cum-laude award outcome, the conclusion of the conditional independence test implies that the data generating mechanism is unfair. Allowing for confounding requires us to use the Bayesian testing procedure proposed in Section 5 for the IV inequalities. We have $|{\cal X}|=|{\cal X}_{D}|=6$, $|{\cal Y}|=|{\cal X}_{A}|=2,|{\cal Z}|=|{\cal X}_{S}|=2$. We choose a flat Dirichlet prior over parameters ($\theta=\left\{P(d,a,s):d\in{\cal X}_{D},a\in{\cal X}_{A},s\in{\cal X}_{S}\right\}$) in both models $\mathbb{M}_{0},\mathbb{M}_{1}$, i.e., for $i=0,1,\pi(\theta|\mathbb{M}_{i})=c_{i}\text{Dir}{\left({1,1,\cdots,1}\right)}$ where $c_{i}$ is a normalizing constant. The counts from the data $R_{1},R_{2}\cdots R_{m}$ are used to obtain the posterior, $P(\theta\mid R_{1},R_{2},\cdots R_{m})$ which is also a Dirichlet distribution. Using $n=10^{6}$ samples, we observe no violations of the IV inequality. Therefore, the confidence interval for the posterior probability of the cum-laude data satisfying the IV inequalities is $\left[1-3.69\times 10^{-6},1\right]$. In contrast to the conclusion one arrives at without allowing for confounding, when allowing for confounding we arrive at a markedly different conclusion that, given the available data and restricting the analysis to only three variables, the fairness of the data-generating mechanism is undecidable.
@@
Frequentist Test of Wang et al. (2017): The frequentist test of Wang et al. (2017) converts every IV inequality into a one-sided association test for a $2\times 2$ contingency table. Specifically, for fixed $d,a$, an IV inequality of the form
@@
Since the direction of the one-sided test matters, we check the sign of the difference of conditional probabilities using maximum likelihood (ML) estimates and then conduct a Pearson’s chi-square test for independence. For the Berkeley data, the ML estimates of $\gamma^{d,a}$ were negative (less than $-0.6$) for all $d,a$, and the independence tests rejected the null hypothesis of independence with p-value $0.0$ (i.e., less than the smallest positive number representable using double precision floating point format, i.e., $<5\times 10^{-324}$). We take this to be significant evidence that the null hypothesis of $\gamma^{d,a}\leq 0$ is not rejected. As noted in Wang et al. (2017), although we test for multiple IV inequalities, the Bonferroni correction is $1/2$ and does not scale as the number of IV inequalities.

### v2

@@ D.1 Sharpness of IV inequalities
For $M\in\mathbb{M}_{\text{IV}}$, the response-function parameterization yields a counterfactually equivalent SCM [Forré and Mooij, 2025, Section 8.4] $\tilde{M}=(V,\tilde{W},\tilde{\mathcal{X}},\tilde{f},\tilde{P})$, where $V=\left\{Z,X,Y\right\},\tilde{W}=\left\{R,U_{Z}\right\},\tilde{\mathcal{X}}=\mathcal{X}_{V}\times\mathcal{X}_{\tilde{W}},\tilde{f}={\left({\tilde{f}_{Z},\tilde{f}_{X},\tilde{f}_{Y}}\right)}$ where we define $\mathcal{X}_{R},\tilde{f},\tilde{P}$ through the function $\Phi:\mathcal{X}_{W}\mapsto\mathcal{X}_{\tilde{W}}$ where
@@
$\mathbb{E}=\left\{\delta_{x,y|0}+\delta_{x^{\prime},y^{\prime}|1}:x,x^{\prime}\in{\cal X};y,y^{\prime}\in{\cal Y}:x\neq x^{\prime}\right\}\cup\left\{\delta_{x,y|0}+\delta_{x,y|1}:x\in{\cal X},y\in{\cal Y}\right\}.$
@@
An elementary result in convex geometry states that a point is an extreme point of a polyhedral set (defined by a set of linear (in)equality constraints) in $\mathbb{R}^{m}$ vector space if and only if it is a feasible point and there exists a subset $A$ of $m$ constraints that are active and linearly independent. For the case at hand, for a point to be an extreme point of $\mathcal{K}_{\text{IV}}$, it has to satisfy the above $26n+$ 2 constraints (feasibility) and additionally, a subset of $n_{X}n_{Y}n_{Z}=4n$ of them must be active and linearly independent. The two normalization constraints are equality constraints and hence must be active at any feasible point.
@@
We first show that all points in $\mathbb{E}$ are extreme points of $\mathcal{K}_{\text{IV}}$. First, choose $x,x^{\prime}\in{\cal X}$ with $x\neq x^{\prime}$ and choose $y,y^{\prime}\in{\cal Y}$. It can be verified that $\delta_{x,y|0}+\delta_{x^{\prime},y^{\prime}|1}$ satisfies the IV inequalities with $2$ out of the $2n$ IV inequalities being active. Further, $24n-$ 2 non-negativity constraints and both normalization constraints are active. For the subset $A$ of active constraints we can take e.g. both active IV inequalities together with the $24n_{X}-$ 2 active nonnegativity constraints; one can check that these are linearly independent. Second, choose $x\in{\cal X},y\in{\cal Y}$. It can be verified that $\delta_{x,y|0}+\delta_{x,y|1}$ satisfies the IV inequalities with $2$ out of the $2n$ IV inequalities being active. Further, $24n-$ 2 non-negativity constraints and both normalization constraints are active. For the subset $A$ of active constraints we can take e.g. both active IV inequalities together with the $24n_{X}-$ 2 active nonnegativity constraints; one can check that these are linearly independent.
@@
Finally, we check whether $\mathbb{E}$ exhausts all extreme points. Pick a feasible point $b\in\mathbb{R}^{n_{X}n_{Y}n_{Z}}$. We will refer to the indices $i$ with $b_{i}\neq 0$ ($b_{i}=0$) as the “non-zero (zero) entries of $b$”, and to the indices $j$ in active constraint $a^{T}b=1$ with $a_{j}\neq 0$ as the “active entries of the constraint”. For a set of entries of $b$ and a set of active entries of one or more active constraints, we refer to their intersection as their “overlap”. We also call the $b_{i}$ corresponding to $K(X,Y|Z=z)$ a “stratum”.
@@
Suppose $b$ is an extreme point of $\mathcal{K}_{\text{IV}}$. Then at least $24n-$ 2 of the $6n$ inequality constraints must be active. Hence if exactly $r$ IV inequalities are active at $b$, then we need at least $4n-2-r$ active nonnegativity constraints at $b$, or in other words, $b$ must have at least $4n-2-r$ zero entries. Because both strata of $b$ need to be normalized, $b$ can have at most $24n-$ 2 zero entries. The number of zero entries in $b$ that can overlap with the active entries of the active IV inequalities is upper bounded by $r$; indeed, otherwise there would be an active IV inequality for which both its active entries are zero entries of $b$, contradicting that this IV inequality is active.
We will show that the only possibilities for $b$ are the extreme points that we already identified, proceeding case by case.
  1. 1.
    $r>2$. This yields a contradiction with the normalization constraints. Indeed, each entry of $b$ appears in exactly one IV inequality, and each IV inequality contains exactly two active entries (one for each stratum). Hence, the sum of those entries of $b$ that correspond with active entries of active IV constraints must be exactly $r$. However, by the normalization constraints, the sum over all entries of $b$ must be 2. This implies that $r\leq 2$.
  2. 2.
    $r=2$. Then $b$ must contain at least $44n-$ 4 zero entries, of which at most two can overlap with the active entries of the active IV inequalities.
    1. (a)
      If there are no such overlaps, then all other entries of $b$ must zero entries. This means that we need the $44n-$ 4 nonnegativity constraints corresponding to those other entries in the subset $A$, as well as the two normalization and inequality constraints; no other active constraints exist that could be added to $A$. But then the constraints in $A$ would not be linearly independent. Indeed, the following linear dependence between the active constraints is obtained: adding the coefficient vectors of the two active IV inequalities and those of all active non-negativity constraints yields the same result as adding the coefficient vectors of the two normalization constraints. So we arrive at a contradiction.
    2. (b)
      If there is at least one overlap, then $b_{j}=1$ with $j$ the other active entry in the active IV inequality with the overlap. This implies $12n-$ 1 zeroes in the stratum of $j$. Consider now the other active IV inequality. Since $b_{j}=1$, the active entry corresponding to that stratum must be a zero entry of $b$. Hence, $b_{k}=1$ for $k$ the other active entry in this IV inequality. This means that $b$ must be of the form $\delta_{x_{1},x_{2}|0}+\delta_{x_{1}^{\prime},x_{2}^{\prime}|1}$, with the two non-zero entries corresponding to active entries of two different IV inequalities, and hence we recover the extreme points already identified.
  3. 3.
    $r=1$. Then $b$ must contain at least $34n-$ 3 zero entries, of which at most one can overlap with the active entries of active IV inequalities.
@@
    Because of the normalization constraints, we need at least one non-zero entry in both strata. This means that there must be a stratum $x_{3}$ with exactly one non-zero entry, and then $b$ must be of the form $b=\delta_{x_{1},x_{2}|x_{3}}+\gamma\delta_{x_{1}^{\prime},x_{2}^{\prime}|x_{3}^{\prime}}+(1-\gamma)\delta_{x_{1}^{\prime\prime},x_{2}^{\prime\prime}|x_{3}^{\prime}}$ with $\gamma\in(0,1]$ and $x_{3}^{\prime}\neq x_{3}$. This means that the active IV inequality must be the one that has active entry $(x_{1},x_{2}|x_{3})$. Its other active entry is then $(x_{1},1-x_{2}|x_{3}^{\prime})$, and this must be an overlapping zero entry of $b$.
    If $\gamma=1$ then this would also activate the IV inequality that contains active entry $(x_{1}^{\prime},x_{2}^{\prime}|x_{3}^{\prime})$. Since the only active IV inequality must also contain active entry $(x_{1},x_{2}|x_{3})$, this gives a contradiction, as the two coefficients of $b$ at the active entries sum to 2 instead of 1. Hence, $\gamma\in(0,1)$, and $b$ contains exactly $34n-$ 3 zero entries.
    So the active constraints consist of 1 active IV inequality, 2 active normalization constraints and $34n-$ 3 active nonnegativity constraints. However, these are not linearly independent. Indeed: subtracting the coefficient vectors of the $12n-$ 1 nonnegativity constraints in stratum $x_{3}$ from the coefficient vector of the normalization constraint of that stratum, and adding the coefficient vector of the nonnegativity constraint corresponding to $(x_{1},1-x_{2}|x_{3}^{\prime})$ gives a vector that is identical to the coefficient vector of the active IV inequality.
    So we have arrived at a contradiction.
  4. 4.
    $r=0$. Then $b$ must contain at least $24n-$ 2 zero entries. Because of the normalization constraints, $b$ needs one non-zero entry in both strata, and its value must be 1. This will activate at least one IV inequality. Contradiction. ∎
@@ D.3 Equivalence of Statistical Tests
Note that ${\cal P}_{\text{IV}+}=\left\{P(Z):\forall z,P(Z=z)>0\right\}\otimes\mathcal{K}_{\text{IV}}$ since assuming positivity, (2) is identical to (3). Further, ${\cal P}_{\text{cf-notion}}=\left\{P_{M}(S):M\in H^{0}_{\text{cf-notion}+}\right\}\otimes{\cal K}_{\text{cf-notion}}$ since assuming positivity, $\mathbb{M}_{\text{cf}+}=\mathbb{M}_{\text{cf}}$ and $P_{M}{\left({D,A\mid\text{do}(S)}\right)}=P_{M}{\left({D,A\mid S}\right)}$ for $M\in\mathbb{M}_{\text{cf}+}$. Since the first factors are identical, Theorem 18 follows from Lemma 26. ∎
@@
D.4 Relaxing the assumption of no confounding between $S$ and $D$
In this section, we prove that Theorem 17 and Theorem 18 hold when $\mathbb{M}_{\text{IV}+}$ and $\mathbb{M}_{\text{cf}+}$ are expanded by allowing for confounding between $S$ and $D$. Denote the corresponding expanded models by $\mathbb{M}_{\text{IV}+ZX}$ and $\mathbb{M}_{\text{cf}+SD}$, respectively, where the former has structural equations of the form $Z=f_{Z}(U_{Z},U_{ZX}),X=f_{X}(Z,U_{X},U_{ZX},U_{XY}),Y=f_{Y}(X,U_{Y},U_{XY})$ where $U_{Z},U_{X},U_{Y},U_{ZX},U_{XY}$ are independent exogenous random variables, and the latter has structural equations of the form $S=f_{S}(U_{S},U_{SD}),D=f_{D}(S,U_{D},U_{SD},U_{DA}),A=f_{A}(S,D,U_{A},U_{DA})$ where $U_{S},U_{D},U_{A},U_{SD},U_{DA}$ are independent exogenous random variables. The corresponding expanded causal null hypothesis corresponding to the fairness notions are denoted by $H^{0}_{\text{cf-notion}+SD}$ where we use ‘notion’ as a placeholder for ‘graph’,‘ctrf’, ‘inter’.
For $M\in\mathbb{M}_{\text{IV}+ZX}$, pick $U\sim\text{Unif}\left[0,1\right]$ and a deterministic map $h:\mathcal{Z}\times\left[0,1\right]\mapsto\mathcal{U}_{ZX}$ such that $U_{ZX}=h(Z,U)$ a.s.. Note that such an $h$ always exists for any random variable $U_{ZX}$ taking values in a standard measurable space (see e.g. [Forré and Mooij, 2025, Corollary 2.7.7]). Define $\tilde{M}=(V=\{\tilde{Z},\tilde{X},\tilde{Y}\},W=\{U_{\tilde{Z}},U_{\tilde{X}},U,U_{\tilde{Y}},U_{\tilde{XY}}\},{\cal X}={\cal X}_{V}\times{\cal X}_{W},\tilde{f}=(f_{\tilde{Z}},f_{\tilde{X}},f_{\tilde{Y}}),\tilde{P})$ where a) $\forall u_{\tilde{Z}},f_{\tilde{Z}}(u_{\tilde{Z}})\triangleq u_{\tilde{Z}}$, b) $\forall u_{\tilde{X}},u,u_{\tilde{XY}},\tilde{z},f_{\tilde{X}}(u_{\tilde{X}},u,u_{\tilde{XY}},\tilde{z})\triangleq f_{X}(\tilde{z},h(\tilde{z},u),u_{\tilde{X}},u_{\tilde{XY}})$, c) $\forall u_{\tilde{Y}},u_{\tilde{XY}},\tilde{x},f_{\tilde{Y}}(u_{\tilde{Y}},u_{\tilde{XY}},\tilde{x})\triangleq f_{Y}(u_{\tilde{Y}},u_{\tilde{XY}},\tilde{x})$, and $\tilde{P}=P_{M}(Z)\otimes P_{X}\otimes\text{Unif}\left[0,1\right]\otimes P_{Y}\otimes P_{XY}$ where $P_{X},P_{Y},P_{XY},P_{Z},P_{ZX}$ are marginals of $P$ over $U_{X},U_{Y},U_{XY},U_{Z},U_{ZX}$ respectively. Note that $P_{M}(Z)$ is the pushforward of $P_{Z}\otimes P_{ZX}$ through $f_{Z}$ thus making $\tilde{P}$ a product distribution over the exogenous random variables. By the above construction, $\tilde{M}\in\mathbb{M}_{\text{IV}+}$. For every $M\in\mathbb{M}_{\text{IV}+ZX}$, $P_{M}(X,Y,Z)=P_{M}(Z)\times P_{M}(X,Y\mid Z)=P_{\tilde{M}}(\tilde{Z})\times P_{\tilde{M}}(\tilde{X},\tilde{Y}\mid\text{do}(\tilde{Z}))\in{\cal P}_{\mathbb{M}_{\text{IV}+}}$. Therefore, $\{P_{M}(Z,X,Y):M\in\mathbb{M}_{\text{IV}+ZX}\}={\cal P}_{\mathbb{M}_{\text{IV}+}}$. Using a similar argument that replaces the labels $Z,X,Y$ by $S,D,A$, $\{P_{M}(S,D,A):M\in H^{0}_{\text{cf-notion}+SD}\}={\cal P}_{\text{cf-notion}}$. This implies Theorem 17 and Theorem 18 hold for $\mathbb{M}_{\text{IV}+ZX}$ and $\mathbb{M}_{\text{cf}+SD}$ respectively.
@@ E.1.1 Counterfactual Fairness and Demographic Parity
We restate the counterfactual notion of fairness from Kusner et al. [2017] for the Berkeley example below.
Definition 27 (Counterfactual Fairness [Kusner et al., 2017]).
@@
The counterfactual fairness notion of Kusner et al. [2017] implies demographic parity for the Berkeley example without allowing for confounding.
@@
Note that demographic parity falls prey to Simpson’s paradox in the Berkeley example. The above result shows that a valid test for demographic parity is a valid test for Kusner et al. [2017]’s counterfactual fairness notion for the assumed model class, $\mathbb{M}_{\text{no-cf}}$.
@@ E.1.2 Path-dependent Counterfactual Fairness
We next show that testing the path-dependent counterfactual fairness notion given in the appendix of Kusner et al. [2017] coincides with a conditional independence test $A\perp\!\!\!\perp S\mid D$.
Definition 29 (Path-dependent Counterfactual Fairness [Kusner et al., 2017]).
@@ E.2 With Confounding
In this section, we compare the statistical tests that result from the NDE that the notions of Nabi and Shpitser [2018] and Chiappa [2019] are based on, and Kusner et al. [2017]’s counterfactual fairness and path-dependent counterfactual fairness notions, when confounding is allowed.
@@ E.2.1 NDE
With confounding between the mediator and the outcome, Kaufman et al. [2005] obtain bounds on the NDE for the all-binary variable case by the linear programming approach of Balke and Pearl [1997]. The resulting bounds are implied by the IV inequalities but not equivalent to them, which gives us a strictly weaker test than the IV inequalities. For completeness, we present the bounds below. The lower and upper bounds for $NDE(A;0\rightarrow 1)$ are
@@
E.2.2 Counterfactual Fairness [Kusner et al., 2017]
@@
E.2.3 Path-dependent Counterfactual Fairness [Kusner et al., 2017]
@@ Appendix F Additional Results for Bayesian Testing Procedure: Cum-laude Dataset and Prior Sensitivity
We consider the dataset from Bol [2023] that contains data from $5239$ PhD students in the Netherlands studying at a large Dutch university from 2011-2021. Bol [2023] observed a bias in the percentage of ‘cum-laude’ distinctions awarded to male PhD students ($6.57\%$) versus female PhD students ($3.68\%$).
@@
As reported by Bol [2023], unlike the Berkeley dataset, the bias among female and male cum-laude award rates does not vanish when conditioned on department. Therefore, with the assumption of no confounding between the academic field choice and the cum-laude award outcome, the conclusion of the conditional independence test implies that the data generating mechanism is unfair when assuming that no latent unprotected mediators between sex and cum-laude award rates exist. Allowing for confounding requires us to use the Bayesian testing procedure proposed in Section 5 for the IV inequalities. We have $|{\cal X}|=|{\cal X}_{D}|=6$, $|{\cal Y}|=|{\cal X}_{A}|=2,|{\cal Z}|=|{\cal X}_{S}|=2$. We choose a flat Dirichlet prior over parameters ($\theta=\left\{P(d,a,s):d\in{\cal X}_{D},a\in{\cal X}_{A},s\in{\cal X}_{S}\right\}$) in both models $\mathbb{M}_{0},\mathbb{M}_{1}$, i.e., for $i=0,1,\pi(\theta|\mathbb{M}_{i})=c_{i}\text{Dir}{\left({1,1,\cdots,1}\right)}$ where $c_{i}$ is a normalizing constant. The counts from the data $R_{1},R_{2}\cdots R_{m}$ are used to obtain the posterior, $P(\theta\mid R_{1},R_{2},\cdots R_{m})$ which is also a Dirichlet distribution. Using $n=10^{6}$ samples, we observe no violations of the IV inequality. Therefore, the confidence interval for the posterior probability of the cum-laude data satisfying the IV inequalities is $\left[1-3.69\times 10^{-6},1\right]$. Hence, when allowing for confounding, on arrives at the conclusion that, given the available data and restricting the analysis to only three variables, the fairness of the data-generating mechanism is undecidable.
@@
Frequentist Test of Wang et al. [2017]: The frequentist test of Wang et al. [2017] converts every IV inequality into a one-sided association test for a $2\times 2$ contingency table. Specifically, for fixed $d,a$, an IV inequality of the form
@@
Since the direction of the one-sided test matters, we check the sign of the difference of conditional probabilities using maximum likelihood (ML) estimates and then conduct a Pearson’s chi-square test for independence. For the Berkeley data, the ML estimates of $\gamma^{d,a}$ were negative (less than $-0.6$) for all $d,a$, and the independence tests rejected the null hypothesis of independence with p-value $0.0$ (i.e., less than the smallest positive number representable using double precision floating point format, i.e., $<5\times 10^{-324}$). We take this to be significant evidence that the null hypothesis of $\gamma^{d,a}\leq 0$ is not rejected. As noted in Wang et al. [2017], although we test for multiple IV inequalities, the Bonferroni correction is $1/2$ and does not scale as the number of IV inequalities.